OpenStudy (anonymous):

The letters r and θ represent polar coordinates. Write the equation using rectangular coordinates (x, y). r = 8cos θ

OpenStudy (anonymous):

When converting to polar, you have \(x=r\cos\theta\), right? How can you get an \(r\) on the right side? Multiply both sides by \(r\): \[r^2=8r\cos\theta\] And what do you know about \(r^2\)?